Circular No. 7612

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ay IN IC 4423 P. Nugent, Lawrence Berkeley National Laboratory (LBNL), on behalf

of G. Aldering, I. Hook, S. Perlmutter, and L. Wang, write: “A spectrum (range 0.5–1.0 μm) obtained on Apr. 22 UT at Keck II (+ ESI) shows that SN 2001ay (cf. IAUC 7611) is a type-Ia supernova, ∼ 1 week prior to maximum light. The redshift of the supernova is consistent with its host galaxy. The Si II feature near 580 nm is pratically absent, while the Si II feature near 615 nm is quite strong (features typical of a normal type-Ia event at an early epoch).”

A. Conley and R. Gibbons, LBNL, report that photometry taken with the Cerro Tololo 0.9-m telescope show that SN 2001ay has brightened since Apr. 18 (see IAUC 7611) and is currently at V = 16.7.

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001ay, obtained by M. Calkins on Apr. 21.47 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), confirms it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9067 km/s for the host galaxy, the supernova expansion velocity is ∼ 14 700 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light. An image taken by Matheson and Jha on Apr. 23.43 with the FLWO 1.2-m telescope (+ 4Shooter) shows the supernova at V = 16.7.

COMET C/1998 V7 (SOHO) Further to IAUCs 7585 and 7606, D. Hammer reports his measure-

ments of another very faint Kreutz sungrazing comet found by M. Oates in archival C2 coronagraph images at the SOHO website. The reduced measurements and orbital elements by B. G. Marsden appear on MPEC 2001-G34.

1998 UT  α_{2000}  δ_{2000}
Nov. 7.221  14^{h}43.^{m}9  −17^{o}34′

SUPERNOVA 2001V IN NGC 3987 CCD R photometry by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): Feb. 24.880 UT, 15.1; 27.837, 14.8; Mar. 6.910, 14.7; 18.851, 15.1; Apr. 1.845, 15.3; 2.843, 15.2; 3.863, 15.4; 14.844, 15.6.

2001 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green